OpenStudy (anonymous):

Solve log½(x - 3) = -4

OpenStudy (anonymous):

@SolomonZelman help me ?

hartnn (hartnn):

use the definition \(\large \log_ab=c \implies b= a^c \) and is the base =1/2 or there is a product of 1/2 and (x-3) ??? \(\log_{1/2} (x-3) \quad or \quad \log (1/2)(x-3) ??\)